If a R200 billion increases in investment spending creates R200 billion of new income in the first round of the multiplier process and R160 billion in the second round. Calculate:
a. the marginal propensity to consume (MPC).
b. the value of the expenditure multiplier in this closed economy.
3.3 Assuming a private closed economy whereby the marginal propensity to consume is 0.9 and investment spending decreases by R1000 billion. What will be the change on equilibrium GDP?
3.4 Suppose that the real GDP increase by R5,000 billion when government expenditure on the construction of new roads increase by R1,500 billion. What is the value of the marginal propensity to consume?
3.5 Assuming that the central government decides to cut taxes by R100 billion to stimulate the economy. The relevant marginal propensity to consume is 0.6 (60 percent). What will be the impact of such fiscal policy on equilibrium GDP?
1
Expert's answer
2020-04-21T18:39:36-0400
  1. a."\Delta Y(1-MPC)=\Delta I"

200(1-MPC)=200

MPC=0

b."M=\frac{\Delta Y}{\Delta I}=\frac{200}{200}=1"

M=ΔI


2 a."\Delta Y(1-MPC)=\Delta I"

160 =(1-MPC)=200

MPC=0.25

b."M=\frac{\Delta Y}{\Delta I}=\frac{160}{200}=0.8"

3.3"\Delta Y(1-MPC)=\Delta I"

Y(1-0.9)=1000

Y=-10 000

3.4

"M=\frac{Y}{G}=\frac{5000}{1500}=3.33"

"M=\frac{1}{1-MPC}"

"3.33=\frac{1}{1-MPC}"

MPC=0.7

3.5"M=\frac{1}{1-MPC}"

M=2.5

"M=\frac{GDR}{T}"

"GDR=\times({2.5}{100}=250"